E(s) = 5. 1. E (21) = 2. E (4r) = 3. B (6-) - 4. E (r) = 5. E(2r + 1) = 6. E (-+ 1) = 7.E (7 - r) =
Answer #1

Answer : E(r) = 5

1. E(2r) = E(2*5) = 10

2. E(4r) = E(4*5) = 20

3. E(1/5r) = E(1/5*5) = 1

4. E(-1/10*r) = E(-1/10*5) = -0.50

5. E(2r + 1) = E(2*5 + 1) = 11

6. E(-1/2+ r) = E(-1/2+5) = 4.5

7. E(7 - 3/5r) = E(7- 3/5*5) = 4